# Extrapolation of Machine-Learning Interatomic Potentials for Organic and Polymeric Systems

### v2

@@ III Results
where $X_{i}$ is the averaged SOAP vector of the atomic environments in the $i$th molecule within the bulk frame. Effectively, this representation re-weights the feature space by reducing (or eliminating) the influence of contributions from intramolecular environments on the SOAP vector. As shown in Fig. 4(b), using this far-sighted representation results in a significant improvement in the number of extrapolative regimes achieved by SOAP-Ridge MLIPs in comparison to the results in Fig. 4(a). Similar trends are observed as in Fig. 2(b): steep improvements in MAEs are observed between ethane and propane, and once again from pentane to hexane. This follows the convergence in chemical environments (Fig. 3, Fig. S4). The results shown in Fig. 4 demonstrate the importance of carefully considering neighbor list construction when making MLIPs for different levels of the energetic hierarchy.
IV Conclusions
In this study, we have demonstrated key behaviors of MLIPs that can be applied to build predictive models for oligomeric systems, with implications on MLIP extrapolation for macromolecular and polymeric materials. First, we demonstrate that, in systems with similar chemistries, extrapolation of forces and energies is possible, provided that those environments present in the target molecule are properly sampled within the smaller analogs. Second, moving from one chemistry to another results in a learnable shift in energies; this can be largely ameliorated by predicting the shift in probability distributions with shifting composition. Finally, we demonstrated a powerful lesson about the hierarchy of energies – intermolecular energetics can be largely extrapolated by shifting the attention in model construction away from nearby neighbors. Given the relative difficulty of computing intra- versus intermolecular energetics, this suggests that the most difficult calculations, those relating the interactions of one molecule to another, are the easiest to translate from one system to another of similar chemistry. These results establish a blueprint for polymeric materials, suggesting that many problems are within reach with careful construction of potential energy models.
V Research Funding
@@
This project was funded by the Wisconsin Alumni Research Fund (RKC) and by NSF through the University of Wisconsin Materials Research Science and Engineering Center (DMR-2309000, AL). NH acknowledges support from the Michael and Virginia Conway Professorship. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
VI Data Availability
@@
The datasets generated and/or analysed during the current study are available in the Zenodo repository at Hooven et al. [51] [DOI: 10.5281/zenodo.17227874].
VII Code Availability
@@
VIII Author Contributions
Project conceptualization (NH, RC), data curation (NH, AL), formal analysis (NH, RC), methodology (NH, RC), supervision (AL, RC), writing (NH, AL, RC).
IX Competing Interests
@@
Appendix A Additional methodological details
MACE MLIPs were constructed using the following command and MACE (v.0.3.13):
Listing 1: Training script for MACE MLIPs
raw_dir=\$(python -c "import mace; print(mace.__file__)"); dir=\${raw_dir:0:-12}
python \$dir/cli/run_train.py \
--name="MACE_model" \
--train_file="\$1" \
--valid_fraction=0.05 \
--config_type_weights=’{"Default":1.0}’ \
--E0s=’{1:-6.492647589968434, 6:-38.054950840332474}’ \
--model="MACE" \
--hidden_irreps=’64x0e + 64x1o’ \
--r_max=5.0 \
--batch_size=1 \
--max_num_epochs=1500 \
--energy_key="energy_eV" \
--forces_key="forces" \
--charges_key="charges" \
--amsgrad \
--restart_latest \
--device=cuda \
--patience=20 \
--default_dtype="float32" \
SOAP vectors for SOAP-Ridge MLIPs were computed using the following hyperparameters with librascal (v.0.0.1):
Listing 2: Hyperparameters of SOAP vectors for SOAP-Ridge MLIPs
"interaction_cutoff": 7,
"max_radial": 8,
"max_angular": 4,
"gaussian_sigma_constant": 0.3,
"gaussian_sigma_type": "Constant",
"cutoff_smooth_width": 0.5,
"radial_basis": "GTO",
"cutoff_function_type": "RadialScaling",
"cutoff_function_parameters": {
"rate": 1.5,
"exponent": 3.0,
"scale": 2.0
Appendix B Supporting figures to the text
(a) Alternative representation of Fig. 2(a).
(b) Alternative representation of Fig. 2(b).
Figure S1: Additional visualizations of MACE models to predict total energy and forces.
Figure S2: Map of the coefficient of determination between true and predicted values (represented as the logarithm of $1-R^{2}$)
(a) Energy-fitting results of MACE MLIP for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(b) Force-fitting results of MACE MLIPs for different train (x-axis) and test (y-axis) pairings, respectively. Color indicates the mean-absolute-error (MAE) in units of meV/atom, respectively.
(c) The relationship between the shift in composition $(x_{C}^{\text{train}}-x_{C}^{\text{test}})$ and the shift in energies for extrapolated MLIPs. The models trained on the methane, ethane, and propane datasets show a divergence from this proportionality, as they are not necessarily considered “well-conditioned.” Marker denotes the testing set, and color denotes the training set.
Figure S3: Analogous to Fig. 2, results of predicting the intramolecular potential energy ($\tilde{E}_{intramolecular}=E_{intramolecular}+E_{atom}$) on n=18n=1-8 alkanes.
(a) Maps of H environments. At a perceptive length of 5Å, we see a convergence of hydrogen environments once reaching pentane.
(b) Maps of CH2 environments. At a perceptive length of 5Å, we see a convergence of CH2 environments once reaching hexane.
(c) Maps of CH3 environments. At a perceptive length of 5Å, we see a convergence of CH3 environments once reaching butane.
Figure S4: Analogous to Fig. 3, maps of different environments for the different training sets, via Principal Components Analysis (left) and Principal Covariates Classification (PCovC[43], right)
(a) Relative errors of SOAP-Ridge potentials trained on $X^{total}$ and intermolecular energies.
(b) Relative errors of SOAP-Ridge potentials trained on $X^{fs}$ and intermolecular energies.
Figure S5: Analogous to Fig. 4, heatmaps of relative errors for different training and testing sets calculated using the mean-absolute-error in meV/atom for testing sets divided by the interpolative error (error of the training alkane on its testing set).

### v3

@@ Results
where $X_{i}$ is the averaged SOAP vector of the atomic environments in the $i$th molecule within the bulk frame. Effectively, this representation re-weights the feature space by reducing (or eliminating) the influence of contributions from intramolecular environments on the SOAP vector. As shown in Fig. 3(b), using this far-sighted representation results in a significant improvement in the number of extrapolative regimes achieved by SOAP-Ridge MLIPs in comparison to the results in Fig. 3(a). Similar trends are observed as in Fig. 1(b): steep improvements in MAEs are observed between ethane and propane/butane, and once again from pentane to hexane. This follows the convergence in chemical environments (Fig. 2, Fig. S4). Furthermore, these results demonstrate that careful consideration of MLIP features can improve performance for targets other than the total potential energy of a system.
Beyond n=1-8 linear alkanes
It is worth inspecting the limitations of these models when considering longer molecules or more complex hydrocarbon architectures. As shown in Fig. 4(a), decane (n=10) and dodecane (n=12), the trends follow those similar to the shorter oligomers within the series: we see a reduction in error when moving from propane to butane (30meV/Å to 5-6meV/Å), and a saturation near 1.5-2 meV/Å  once training on hexane through octane. This supports our earlier observations considering the saturation of interaction environments, and suggests that these trends should continue to longer linear alkanes. Returning to the SOAP-Ridge MLIPs constructed using total SOAP vectors ($X^{total}$) and far-sighted SOAP vectors ($X^{fs}$) in Fig. 4(c), we also maintain previous trends: intermolecular energy MAEs decrease consistently as the training alkane chain length increases, and MLIPs constructed using $X^{fs}$ consistently outperform MLIPs constructed with $X^{total}$. This re-confirms our earlier claim that re-weighting the feature-space to exclude intramolecular contributions enables higher accuracy when targeting intermolecular energy.
Molecules of different architectures fare less well in extrapolation. For cyclohexane, 4-propylheptane, and 3,3-diethylpentane, we typically converge to higher values (6-10meV/Å) than within the linear series (1-2meV/Å). On one hand, this is due to the lack of longer alkyl chains within cyclohexane, 4-propylheptane, and 3,3-diethylpentane – moving to hexane, heptane, and octane training data is not nearly as beneficial when the longest uninterrupted alkyl chain is three carbons long. For 4-propylheptane and 3,3-diethylpentane, the larger errors are largely attributed to the tertiary and quaternary carbons (Fig. 4(b)), neither of which exists in a linear alkane. Though we see higher errors from the MACE predicted forces for these molecules, we maintain consistent trends in intermolecular energies as predicted by SOAP-Ridge MLIPs with decreasing errors as training alkane chain length increases, and lower errors for all but one far-sighted MLIPs.
Cyclohexane presents a more interesting case. The carbons in cyclohexane have long received different parameterization than those in hexane via popular force fields such as OPLS-AA[34, 18, 32]. This has largely been attributed to the constrained torsion angles in cyclohexane, which remove the freely rotating C-C dihedrals that characterize linear alkanes[34]. However, here, the key difference is the local environment distribution: linear alkanes provide very few CH2 environments with neighboring CH2 at the 232-3Å range, whereas cyclohexane contains several such neighbors because of its compact ring geometry. As a result, when the model trained only on octane is evaluated on cyclohexane, it encounters CH2–CH2 configurations that were poorly sampled in the training set, producing the shift observed in Fig. 4(b).
Conclusions
In this study, we have demonstrated key behaviors of MLIPs that can be applied to build predictive models for oligomeric systems, with implications on MLIP extrapolation for macromolecular and polymeric materials. First, we demonstrate and visualize that, in systems with similar chemistries, extrapolation of forces and energies is possible, provided that those environments present in the target molecule are properly sampled within the smaller analogs. Specifically for alkanes, we see a strong reduction in error with butane, and we see the force errors converging once we reach hexane. These results were also demonstrated on longer linear alkanes and different architectures. The latter proved more nuanced and complicated as an extrapolation task, as while non-linear alkanes contain the same chemistry, MLIPs do not consider these aspects explicitly, but rather as emerging from data distributions. And, because the environmental distributions of branched or cyclical alkanes are fundamentally different from linear alkanes, the MLIPs fare worse in extrapolating to these systems.
Second, we showed that extrapolation of energies results in a learnable shift between different molecules; this can be largely ameliorated by predicting the mean shift in probability distributions as a function of composition. As forces are not affected by constant shifts in energy, the extrapolation of forces between different molecules is straightforward and does not require accounting for this effect.
Finally, we demonstrated a powerful lesson about the hierarchy of energies. Intermolecular contributions are important in determining thermodynamic properties, but have historically been challenging to predict due to their small magnitude relative to the intramolecular component. We have shown that careful partitioning of molecular descriptors can yield MLIPs capable of predicting and extrapolating intermolecular energetics. Given the relative difficulty of computing intra- versus intermolecular energetics, this suggests that the most difficult calculations, those relating the interactions of one molecule to another, are the easiest to translate from one system to another of similar chemistry.
Research Funding
@@ Conclusions
This project was funded by the Wisconsin Alumni Research Fund (RKC) and by NSF through the University of Wisconsin Materials Research Science and Engineering Center (DMR-2309000, AL, CC). NH acknowledges support from the Michael and Virginia Conway Professorship. The funders played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
Data Availability
@@
The datasets generated and/or analysed during the current study are available in the Zenodo repository at [27] [DOI: 10.5281/zenodo.17227874].
Code Availability
@@
Author Contributions
Project conceptualization (NH, RC), data curation (NH, CC, AL), formal analysis (NH, RC), methodology (NH, RC), supervision (AL, RC), writing (NH, AL, CC, RC).
Competing Interests